OpenStudy (anonymous):

#2. 3(y – 3) = 2y – 9 + y (1 point) a; one solution b; no solutions c ;infinitely many solutions

myininaya (myininaya):

Simply both sides first

myininaya (myininaya):

What is 3(y-3)=? (hint: use distributive property) What is 2y-9+y=? (hint: combine like terms)

OpenStudy (anonymous):

i dunno how to do that . XD

OpenStudy (anonymous):

im stupid . CX

OpenStudy (anonymous):

just . say a b or c ♥

myininaya (myininaya):

Distributive property a(b+c)=ab+ac Like terms are terms that have the same variable part. If I asked you to simplify x+x you would say 1x+1x=2x Or if you had x+2x you would say 1x+2x=3x You aren't stupid and you can do this.

OpenStudy (anonymous):

asdfghjkl . ;o

myininaya (myininaya):

Come on. Work with me.

myininaya (myininaya):

3(y-3) a(b+c)=ab+ac Or you might like to use this one more: a(b-c)=ab-ac since you have 3(y-3)

myininaya (myininaya):

To each of the terms inside the parenthesis you will multiply those terms by 3

myininaya (myininaya):

So 3(y-3)=3*y-3*3 Correct? Can you simplify this more?

myininaya (myininaya):

You know how to multiply 3 and 3 so the answer is yes. Please write what the left hand side of the equation is simplified.

myininaya (myininaya):

And then we will look at the right together.

OpenStudy (anonymous):

oh my god . this is too complicated for me . is it a ; b ; or c ;

myininaya (myininaya):

@hurrrdawg This is not the site where you will find just answers. You have to be willing to learn how to get them here. It is against the code of conduct for you to just be looking for answers.

myininaya (myininaya):

Perhaps, you would like to read it: http://openstudy.com/code-of-conduct

myininaya (myininaya):

It is not too complicated. You can do this. Don't just give up already. That will not help you.

OpenStudy (anonymous):

@myininaya I'm not looking for the answer or anything, but could you explain the answer. Doesn't y get cancelled? (or am I being dumb)